# Submitted to Astronomy Letters THE ERUPTIVE YOUNG STAR IRAS 21204+4913

### v1

@@ 3.2 Photometry
Fig. 3 shows the evolution of the brightness of IRAS 21204+4913 in the optical and NIR bands during our observation period, along with several of its color indices. It is evident that the rapid brightening in the optical (by nearly during November 2025; Stecklum 2025) slowed down significantly in December and essentially ceased by January. Throughout our monitoring period, the source’s brightness variations from 0.37 to 4.8 m occurred nearly simultaneously. More precisely, after December 10, 2025, the color indices and remained constant within the photometric uncertainties, while the color initially decreased slightly and then stabilized.
@@ 3.4 Spectra
Nevertheless, the spectrum of IRAS 21204+4913 exhibits features that are atypical for FUors. First, as seen in Fig. 5, the TiO molecular bands appear in emission rather than absorption. Second, there is a clear dependence both of the line width and radial velocity on the excitation energy (Fig. 7) – a behavior not previously observed in classical FUors, to the best of our knowledge.
@@ 4.5 Accretion rate estimate
Fig. 12 shows the SED of IRAS 21204+4913 in the m range, uncorrected for extinction. The SED is constructed from photometric data obtained at the beginning (December 9, 2025) and the end (January 29, 2026) of our observations (see Tables A1 and A2). During this period, the total flux in the considered wavelength range increased by approximately %, reaching a value of erg s-1cm-2. At a distance of 500 pc, this corresponds to a luminosity of L
@@ 5 Conclusion
We also discovered for the first time an emission line at Å and TiO molecular emission bands in the spectrum of an eruptive YSO with an absorption-dominated continuum. It is difficult to assess how critical these emission features are for understanding the outburst mechanism of IRAS 21204+4913. However, we note that the CO bands near m – typically seen in absorption in classical FUors – are observed in emission in the eruptive YSOs V1647 Ori and IRAS 06297+1021W (Connelley and Reipurth, 2018), which led Contreras Peña,Carlos et al. (2025) to classify them as intermediate objects between FUors and EXors.

### v2

# Accepted by Astronomy Letters THE ERUPTIVE YOUNG STAR IRAS 21204+4913

@@ 3.2 Photometry
Fig. 3 shows the evolution of the brightness of IRAS 21204+4913 in the optical and NIR bands during our observation period, along with several of its color indices. It is evident that the rapid brightening in the optical (by nearly during November 2025; Stecklum 2025) slowed down significantly in December. The stellar brightness reached maximum in January and then began to gradually decrease in optical band. Throughout our monitoring period, the source’s brightness variations from 0.37 to 4.8 m occurred nearly simultaneously. More precisely, after December 10, 2025, the color indices and remained constant within the photometric uncertainties, while the color initially decreased slightly and then stabilized.
@@ 3.4 Spectra
Nevertheless, the spectrum of IRAS 21204+4913 exhibits features that are atypical for FUors. First, as seen in Fig. 5, the TiO molecular bands appear in emission rather than absorption that has already been observed in the young eruptive stars LkH 225S and V2429 Cyg as well as in IRAS 05451+0037 classified as Class I/II YSO (Hillenbrand et al., 2012, 2022). Second, there is a clear dependence both of the line width and radial velocity on the excitation energy (Fig. 7) – a behavior not previously observed in classical FUors, to the best of our knowledge.
@@ 4.5 Accretion rate estimate
Fig. 12 shows the SED of IRAS 21204+4913 in the m range, uncorrected for extinction. The SED is constructed from photometric data obtained at the beginning (December 9, 2025) and the middle (January 29, 2026) of our observations (see Tables A1 and A2). During this period, the total flux in the considered wavelength range increased by approximately %, reaching a value of erg s-1cm-2. At a distance of 500 pc, this corresponds to a luminosity of L
@@ 5 Conclusion
We also discovered for the first time an emission line at Å and TiO molecular emission bands in the spectrum of an eruptive YSO with an absorption-dominated continuum. It is difficult to assess how critical these emission features are for understanding the outburst mechanism of IRAS 21204+4913 and other young eruptive stars. However, we note that the CO bands near m – typically seen in absorption in classical FUors – are observed in emission in the eruptive YSOs V1647 Ori and IRAS 06297+1021W (Connelley and Reipurth, 2018), which led Contreras Peña,Carlos et al. (2025) to classify them as intermediate objects between FUors and EXors.